### v1

@@ Abstract
We investigate the optimal charging strategy for large-scale electric vehicles in smart grids using a finite-horizon framework within a mean field game approach. By proving the existence and uniqueness of the solution to the consistency condition equation, we analyze the optimal charging strategies for electric vehicles in the mean field limit case. We then construct the approximate optimal charging strategies for a finite number of vehicles in both non-cooperative and cooperative games, and give a numerical experiment accordingly. Lastly, we present numerical analyses to illustrate approximate strategies of non-cooperate and cooperate game.
@@ 1 Introduction
In the context of large-scale network structures, mean field games (MFGs) were independently proposed by Huang et al. (2006) and Lasry and Lions (2007), which devoted to the analysis of dynamic systems where a large number of players interact strategically with each other. By employing the “mean field” approximation, MFGs simplify the complex multi-player game problem into an interaction between a single player and the aggregate behavior of the population. This framework provides an effective analytical tool for addressing the interactions among players in large-scale systems. For more comprehensive insights into MFG theory, the reader may refer to Bensoussan et al. (2016), Carmona and Delarue (2018), Gomes and Saúde (2014) and the additional references cited within those works. Its application has spanned a wide range of fields including economics and finance  (Lacker and Zariphopoulou, 2019; Nuno and Moll, 2018), large networks or graphs (Lacker and Soret, 2022), operations research (Wang and Huang, 2019; Wang et al., 2022), engineering and machine learning (Zhou and Xu (2022)), epidemics control(Laguzet and Turinici (2015); Lee et al. (2021)), smart grid (Gomes and Saúde (2021)) and more. For the tractable case, the MFG method is also intensively studied in the LQG framework, due to the elegance of its analytical tractability. In particular, Huang (2010) introduces LQG games that feature a major player with substantial influence and a multitude of minor players. Subsequently, Bensoussan et al. (2016) provide a comprehensive study of a general class of MFGs in the LQ framework. Zhang and Li (2019) explore linear-quadratic MFGs described by forward-backward stochastic differential equations. Gomes et al. (2023) study the case of linear dynamics for the supply and a cost function consisting of price, control and commodity value, which extends the LQG framework by introducing the product of control and price in the cost function. MFGs offer a natural and elegant mathematical approach to these challenges, delivering decentralized, scalable, and theoretically grounded solutions for optimizing the performance of smart grids.
@@
Social optimum solutions for MFG are also well documented. For example, Li et al. (2016) examine the relationship between mean-field games and social welfare optimization problems. Salhab et al. (2016) focus on dynamic collective choice problems, and address scenarios where a large number of players cooperatively choose between multiple destinations while influenced by group behavior. Feng et al. (2021) involve a major agent and numerous minor agents and investigates a mixed stochastic LQG social optimization. Moreover, Wang and Huang (2019) consider both the non-cooperative and cooperative solutions for a system with sticky prices and adjustment costs. It proposes an auxiliary limiting optimal control problem subject to consistent mean field approximations, and the method is enlightening to our work.
@@
Moreover, versatile framework for modeling price dynamics in markets with numerous interacting agents are applied in MFGs. Notable approaches are listed below: Gomes and Saúde (2021) introduce a price formation model where numerous small players can store and trade a commodity like electricity; Gomes et al. (2020) propose a MFG model for price formation of a commodity with production subject to random fluctuations. The dynamic game models have been extensively utilized to analyze duopolistic competition with sticky prices. Existing literature related to it includes Cellini and Lambertini (2004), Wiszniewska-Matyszkiel et al. (2015), Valentini et al. (2021), Kańska and Wiszniewska-Matyszkiel (2022) and Hoof (2021). Besides, Cellini and Lambertini (2004) and Wiszniewska-Matyszkiel et al. (2015) propose an evolution law for the sticky price in dynamic oligopoly, which will be applied to our work.
@@
This paper aims to develop a strategic framework for a large market composed of numerous agents, using the methodology of MFG. Within a setting characterized by sticky prices and finite time horizons, we investigate how agents can optimize their behavior to minimize expected losses. In the cooperative game case, our objective is to reduce the average social cost across the market. On the other hand, under the non-cooperative framework, we focus on constructing an approximate Nash equilibrium that captures the decentralized decision-making of individual agents. This dual perspective allows for a comprehensive analysis of collective and competitive behaviors in large-scale smart grid systems. Given the close resemblance between our model and the models of Gomes et al. (2023) and Wang and Huang (2019), we now emphasize the distinctions in both model structure and methodological approach. The model in Gomes et al. (2023) determines the mean process of the control variable through a stochastic differential equation. Subsequently, by employing variational methods, they seek a price process such that, when agents take optimal actions to minimize transaction costs, the market clears and supply meets demand. Unlike in their work, our price process is endogenously generated within the system and is a stochastic process driven by control variables. The control variables can be freely adjusted, and the agents’ commodities are subject to independent noise disturbances. The more fundamental difference lies in our objective: we aim to construct $\epsilon$-optimal solutions for both cooperative and non-cooperative games in a large system, rather than achieving supply-demand equilibrium. Besides, our methodological approach mainly involves approximating stochastic processes using fixed functions, rather than employing variational techniques. Different from Wang and Huang (2019), which aims to find an $\epsilon$-Nash equilibrium and $\epsilon$-socially optimal solution for the system based on cost minimization within an infinite-horizon framework, we consider a finite-horizon setting and adopt a cost function of the LQG form, which deduces to the feedback strategies that depend on the state process. More importantly, our cost function is positively correlated with the product of price and demand, while this term is negatively correlated with the cost function proposed by Wang and Huang (2019). This positive correlation introduces great amount of complexity to the problem, as it implies that increasing the norm of the control variable can significantly reduce one component of the cost function to a very small negative number, potentially undermining the coercivity condition. Overcoming the complications introduced by this correlation constitutes the principal difficulty of the present work and serves as the guiding objective behind many of the subsequent proofs.
@@
Our mathematical contributions can be summarized in three aspects: First, within a finite-horizon framework, we propose a mean-field interaction cost function, through which we derive approximate solutions for both the cooperative and non-cooperative games. To the best of our knowledge, it is relatively rare in the literature to address both cooperative and non-cooperative formulations simultaneously in a finite-horizon framework. Second, for specific matrix equations, we construct the solution induced by specific initial values and study the existence and uniqueness of the mean field through its properties. Most of the existing literature employs the Banach fixed-point theorem or other methods to find equilibrium points. In contrast to these approaches, our method is able to more precisely characterize the state of the mean field. Third, we apply the Positive Real Lemma to demonstrate that certain parameter restrictions can ensure that the coercivity condition holds, thereby effectively addressing the challenge posed by the positive correlation between the cost function and the product of price and demand term. Although Wang and Huang (2019) also employs the Positive Real Lemma, the problem we consider is more complex, and the manner in which we utilize the lemma is correspondingly more intricate.
@@
The rest of the paper is organized as follows: In Section 2, we introduce the model of the smart grid, which contains a large number of electric vehicles. We then construct the $\epsilon$-Nash equilibrium charging strategies for the EVs in the non-cooperate games. Afterwards, we construct the asymptotic social optimum charging strategies for the cooperate games. In Section 3, we give numerical simulations for the charging strategies proposed in both the non-cooperate games and cooperate games. Section 4 concludes.
@@ 2 The Model and the Optimal Strategies
In this section, we propose the sticky price model and the optimal charging strategies of smart grid with numerous agents (e.g. EVs). First, we introduce our model. Let the number of agents in the smart grid be $N$. The agents control their electricity (also referred to as commodities) through controlling the charging rate over a shared common time horizon $[0,T]$. Let $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ be a complete filtered probability space with the filtration $\mathbb{F}=(\mathcal{F}_{t})_{t\in[0,T]}$ satisfying the usual conditions, which supports the independent $N$-dimension standard Brownian motions $(W_{t}^{1},\cdots,W_{t}^{N})_{t\in[0,T]}$. Denote the electricity held by agent $i$ at time $t$ by $X_{t}^{i}$. The dynamic of $X^{i}$ satisfies
@@
Here, the parameter $\sigma_{i}\geq 0$ measures the volatility of agent $i$’s electricity, and the stochastic noise term represents uncertainty or variability in the smart grid system, such as grid disturbances or physical noise, communication or control delays, or user behavior uncertainty. In addition, we assume that agent $i$ is equipped with an initial electricity $x_{0}^{i}\in\mathbb{R}$, As shown in (2.1), the control variable for agent $i$ is denoted by $v^{i}$, also referred to as trading rate (or charging rate).
@@
where $\alpha,\beta>0$ are constant market factors, and the initial price is $p_{0}$. The parameter $\alpha$ measures the sensibility of $P$ to current situations. $Q_{t}$ is the average trading rate at time $t$. In this paper, it is equal to $\frac{1}{N}\sum_{i=1}^{N}v_{t}^{i}$.
@@
respectively. The parameter $\zeta$ corresponds to the preferred final storage, and $\kappa$ is the preferred instantaneous storage. We make the assumption that $\eta>0,\ \gamma\geq 0,\ c>0,\ \kappa\in\mathbb{R},\ \text{and}\ \zeta\in\mathbb{R}$ are constants. To introduce the admissible strategy set of $v^{i}$, we assign every stochastic control $w$ the norm $\left\|w\right\|=\sqrt{\mathbb{E}\left[\int_{0}^{T}w_{t}^{2}dt\right]}$. Then, $v^{i}$ should belong to the admissible strategy set
[…]
@@
A set of admissive strategies $\hat{\boldsymbol{v}}=\{\hat{v}^{1},\cdots,\hat{v}^{N}\}\in\mathcal{A}^{N}$ is an $\epsilon$-Nash equilibrium if
$J_{i}(\hat{v}^{i},\hat{\boldsymbol{v}}^{-i})-\epsilon_{N}\leq\inf_{v^{i}\in\mathcal{A}}J(v^{i},\hat{\boldsymbol{v}}^{-i})\leq J_{i}(\hat{v}^{i},\hat{\boldsymbol{v}}^{-i}),\quad i=1,\ldots,N,$ (2.4)
@@
In order to further characterize the rate of $\epsilon$ reduction, we introduce the following notion. A number dependent on $N$ is called an $O(\frac{1}{\sqrt{N}})$ number if there exists a fixed constant $M$ (depending on all the parameter introduced, including $p_{0}$, $T$, and others), which is independent of $N$, such that for $N$ large enough (i.e., there exists a constant $N_{0}>0$, such that for $\forall N>N_{0}$), this number’s absolute value does not exceed $\frac{M}{\sqrt{N}}$. Similarly, we can define $O(1)$ numbers and $O(\frac{1}{N})$ numbers. Apart from that, when we state that a certain quantity dependent on $N$ (for example, a function or a variable $h(N)$) satisfies $h(N)\leq O(\frac{1}{\sqrt{N}})$, we mean that there exists an $O(\frac{1}{\sqrt{N}})$ number, such that $h(N)$ is always no greater than this $O(\frac{1}{\sqrt{N}})$ number. We introduce a similar form of comparing rule for $O(1)$ and $O(\frac{1}{N})$ numbers in a completely analogous way.
Under the second case (namely, the cooperate game case), instead of controlling the commodities independently, the agents cooperate with each other, and aim to achieve the best collective outcome. In this article, the collective outcome represents the average value of $J_{i}(v^{i},\boldsymbol{v}^{-i})$. It reads
@@
$|J_{soc}(\check{v}^{1},\check{v}^{2},\cdots,\check{v}^{N})-\inf_{v^{i}\in\mathcal{A},1\leq i\leq N}J_{soc}(v^{1},v^{2},\cdots,v^{N})|\leq\epsilon_{N},$ (2.5)
where $\epsilon_{N}$ must vanish to 0 as $N\to\infty$.
The restriction for $\epsilon_{N}$ reflects that the expected cost under the proposed strategies converges to that of the best strategies as $N$ goes to $\infty$.
@@
One more point needs to be stated in advance. In this paper, when we solve first-order ordinary differential equations, unless otherwise specified, we are looking for classical solutions (i.e., continuously differentiable solutions), not weak or viscosity ones.
2.1 Non-cooperate Game
In this subsection, we aim to find an $\epsilon$-Nash equilibrium for this system. We will use a mean-field approximation approach, in which we construct $C[0,T]$ functions to approximate certain processes. Our approach can be outlined in the following steps.
@@ 2.1 Non-cooperate Game
    The first step in constructing a mean-field approximation is to approximate $Q$ with a continuous function $\bar{Q}=(\bar{Q}(t))_{t\in[0,T]}$. It follows from (2.2) that the approximate price process $(\bar{P}(t))_{t\in[0,T]}$ satisfies
@@
    For each agent $i$, solve the auxiliary control problem (2.7) to find the optimal control $\hat{v}^{i}$. Calculate the corresponding average optimal control and remove the noise term to obtain a determinist process. By requiring that this process equal $\bar{Q}$, we determine the exact form of $\bar{Q}$.
@@
The rationale for proposing the first two steps is primarily based on the following reasons. Firstly, when the population enlarges to $\infty$, the representative agent has no influence on $Q$, and therefore has no influence on $P$, as but one agent amid a continuum. From its perspective, the process $P$ should be evolved in its cost as a deterministic function. Secondly, the optimization problem for (2.3) becomes much easier to solve once we treat $Q$ and $P$ as fixed continuous functions, instead of random ones. For each $\bar{Q}$, a unique process $\bar{P}$ is determined by (2.6). Each agent determines its optimal control and gives the corresponding average investment process according to $\bar{P}$. If the expectation of this average investment process is equal to $\bar{Q}$, a closed-loop is formed. Within this loop, the strategy adopted by each agent yields outcomes that closely approximate those of the actual optimal strategy as the number of agents increases.
For the given process $\bar{Q}\in C[0,T]$, the approximate price process $\bar{P}$ is uniquely determined by (2.6). We propose the value function of the auxiliary control problem (2.7),
$K_{i}(t,x)=\inf_{v^{i}\in\mathcal{A}}\mathbb{E}\left[\int_{t}^{T}(L(X_{s}^{i},v_{s}^{i})+\bar{P}(s)v_{s}^{i})ds+\Psi(X_{T}^{i})\right].$
where $t$ represents the starting time, and $x$ represents the initial value of $X^{i}$. By dynamic programming principle, the HJB equation satisfies
$\frac{\partial{K}_{i}(t,x)}{\partial t}+\inf_{v^{i}\in\mathcal{A}}\left\{v_{t}^{i}\frac{\partial{K}_{i}(t,x)}{\partial x}+\frac{c}{2}({v_{t}^{i}})^{2}+v_{t}^{i}\bar{P}(t)\right\}+\frac{\sigma_{i}^{2}}{2}\frac{\partial^{2}{K}_{i}(t,x)}{\partial x^{2}}+\frac{\eta}{2}(x-\kappa)^{2}=0,$ (2.8)
along with the terminal condition $K_{i}(T,x)=\frac{\gamma}{2}(x-\zeta)^{2}$. We make the ansatz ${K}_{i}(t,x)=a(t)x^{2}+B(t)x+F_{i}(t)$ yields. Furthermore, the optimal feedback control achieving the minimum is given by
$\hat{v}_{t}^{i}=-\frac{\bar{P}(t)+2a(t)x+B(t)}{c},\quad\forall t\in[0,T].$ (2.9)
Plug (2.9) into (2.8) to obtain the following ODE system:
$\qquad\qquad\qquad a^{\prime}(t)+\frac{\eta}{2}-\frac{(2a(t))^{2}}{2c}=0,$ $a(T)=\frac{\gamma}{2},$ (2.10)
$\qquad\qquad\qquad B^{\prime}(t)-\eta\kappa-\frac{2a(t)(\bar{P}(t)+B(t))}{c}=0,$ $B(T)=-\gamma\zeta,$ (2.11)
$\qquad\qquad\qquad F^{\prime}_{i}(t)-\frac{(\bar{P}(t)+B(t))^{2}}{2c}+\frac{\eta\kappa^{2}}{2}+\sigma_{i}^{2}a(t)=0,$ $F_{i}(T)=\frac{\gamma\zeta^{2}}{2}.$ (2.12)
To make the ansatz hold, we will verify that (2.10),(2.11) and (2.12) give a solution. We thereby introduce subsection 2.1.
@@
The solution for (2.10), (2.11) and (2.12) exists and is unique.
The proofs of the Propositions in this paper have been left to Appendix B. We next present the verification theorem on the best response control for minimizing (2.7), which demonstrates that our ansatz holds.
@@
$\hat{v}_{t}^{i}(t,X_{t}^{i})=-\frac{\bar{P}(t)+2a(t){X}_{t}^{i}+B(t)}{c},\quad\forall t\in[0,T].$
Then, the SDE
@@
admits a unique solution, denoted by $\hat{X}_{t}^{i}$, and $\hat{v}^{i}$ is an optimal Markovian control for minimizing (2.7).

### v2

@@ Abstract
We study the optimal charging strategies for large-scale electric vehicles in smart grids within a finite-horizon mean field game framework. We first establish the existence and uniqueness of the solution to the consistency condition equation, which characterizes the optimal charging behavior in the mean field limit. Building on this result, we construct approximate optimal charging strategies for a finite population of vehicles in both non-cooperative and cooperative settings. Finally, we provide numerical analyses that illustrate and compare the approximate strategies in the non-cooperative and cooperative games.
@@ 1 Introduction
In the context of large-scale network structures, mean field games (MFGs) were independently proposed by Huang et al. (2006) and Lasry and Lions (2007), which devoted to the analysis of dynamic systems where a large number of players interact strategically with each other. By employing the “mean field” approximation, MFGs simplify the complex multi-player game problem into an interaction between a single player and the aggregate behavior of the population. This framework provides an effective analytical tool for addressing the interactions among players in large-scale systems. For more comprehensive insights into MFG theory, the reader may refer to Bensoussan et al. (2016), Carmona and Delarue (2018), Gomes and Saúde (2014) and the additional references cited within those works. Its application has spanned a wide range of fields including economics and finance  (Lacker and Zariphopoulou, 2019; Nuno and Moll, 2018), large networks or graphs (Lacker and Soret, 2022), operations research (Wang and Huang, 2019; Wang et al., 2022), engineering and machine learning (Zhou and Xu (2022)), epidemics control (Laguzet and Turinici (2015); Lee et al. (2021)), smart grid (Gomes and Saúde (2021)) and more.
@@
Social optimum solutions for MFG are also well documented. For example, Li et al. (2016) examine the relationship between mean-field games and social welfare optimization problems. Salhab et al. (2016) focus on dynamic collective choice problems, and address scenarios where a large number of players cooperatively choose between multiple destinations while influenced by group behavior. Feng et al. (2021) involve a major agent and numerous minor agents and investigates a mixed stochastic LQG social optimization.
@@
Moreover, versatile framework for modeling price dynamics in markets with numerous interacting agents are applied in MFGs. Notable approaches are listed below: Gomes and Saúde (2021) introduce a price formation model where numerous small players can store and trade a commodity like electricity; Gomes et al. (2020) propose a MFG model for price formation of a commodity with production subject to random fluctuations. The dynamic game models have been extensively utilized to analyze duopolistic competition with sticky prices. Existing literature related to it includes Cellini and Lambertini (2004), Wiszniewska-Matyszkiel et al. (2015), Valentini et al. (2021), Kańska and Wiszniewska-Matyszkiel (2022) and Hoof (2021).
This paper aims to develop a strategic framework for a large market composed of numerous agents, using the methodology of MFG. Within a setting characterized by sticky prices and finite time horizons, we investigate how agents can optimize their behavior to minimize expected losses. In the cooperative game case, our objective is to reduce the average social cost across the market. On the other hand, under the non-cooperative framework, we focus on constructing an approximate Nash equilibrium that captures the decentralized decision-making of individual agents. This dual perspective allows for a comprehensive analysis of collective and competitive behaviors in large-scale smart grid systems.
@@
Given the close resemblance between our model and the models of Gomes et al. (2023), we now emphasize the distinctions in both model structure and methodological approach. The model in Gomes et al. (2023) determines the mean process of the control variable through a stochastic differential equation. Subsequently, by employing variational methods, they seek a price process such that, when agents take optimal actions to minimize transaction costs, the market clears and supply meets demand. Unlike in their work, our price process is endogenously generated within the system and is a stochastic process driven by control variables. The control variables can be freely adjusted, and the agents’ commodities are subject to independent noise disturbances. The more fundamental difference lies in our objective: we aim to construct $\epsilon$-optimal solutions for both cooperative and non-cooperative games in a large system, rather than achieving supply-demand equilibrium. These distinctions lead us to adopt a completely different methodological approach in our analysis– our methodological approach mainly involves the standard MFE approach, that is, to approximate stochastic processes using fixed functions, rather than employing variational techniques.
@@
Our mathematical contributions can be summarized in three aspects: First, within a finite-horizon framework, we propose a mean-field interaction cost function, through which we derive approximate solutions for both the cooperative and non-cooperative games, while the literature contains relatively few works that offer both approximate Nash equilibria and approximate social optima. Second, for specific matrix equations, we construct the solution induced by specific initial values and study the existence and uniqueness of the mean field through its properties. Most of the existing literature employs the Banach fixed-point theorem or other methods to find equilibrium points. In contrast to these approaches, our method is able to more precisely characterize the state of the mean field. Third, we apply the Positive Real Lemma to demonstrate that certain parameter restrictions can ensure that the coercivity condition holds, thereby effectively addressing the challenge posed by the positive correlation between the cost function and the product of price and demand term. To the best of our knowledge, this challenge has not been addressed in the existing literature.
@@
The rest of the paper is organized as follows: In Section 2, we introduce the model of the smart grid, which consists of a large number of electric vehicles. In Section 3, we construct the $\epsilon$-Nash equilibrium charging strategies for the EVs in the non-cooperate games. Afterwards, in section 4, we construct the asymptotic social optimum charging strategies for the cooperate games. In Section 5, we give numerical simulations for the charging strategies proposed in both the non-cooperate games and cooperate games. Section 6 concludes.
@@ 2 The Model and the Optimal Strategies
In this section, we propose the sticky price model and the optimal charging strategies of smart grid with numerous agents (e.g. EVs). First, we introduce our model. Let the number of agents in the smart grid be $N$. The agents control their electricity (also referred to as commodities) through controlling the charging rate over a shared common time horizon $[0,T]$. Let $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ be a complete filtered probability space with the filtration $\mathbb{F}=(\mathcal{F}_{t})_{t\in[0,T]}$ satisfying the usual conditions, which supports the independent $N$-dimension standard Brownian motions $(W_{t}^{1},\cdots,W_{t}^{N})_{t\in[0,T]}$. Denote the electricity held by agent $i$ at time $t$ by $X_{t}^{i}$. Let the dynamics of $X^{i}$ satisfy
@@
Here, the parameter $\sigma_{i}\geq 0$ measures the volatility of agent $i$’s electricity, and the stochastic noise term represents uncertainty or variability in the smart grid system, such as grid disturbances or physical noise, communication or control delays, or user behavior uncertainty. In addition, we assume that agent $i$ is equipped with an initial electricity $x_{0}^{i}\in\mathbb{R}$, As in (2.1), the control variable for agent $i$, also referred to as trading rate (or charging rate) is denoted by $v^{i}$. To introduce the admissible strategy set of $v^{i}$, we assign every stochastic control $w$ the norm $\left\|w\right\|=\sqrt{\mathbb{E}\left[\int_{0}^{T}w_{t}^{2}dt\right]}$. Then, $v^{i}$ should belong to the admissible strategy set
[…]
@@
where $\alpha,\beta>0$ are constant market factors, and the initial price is $p_{0}$. The parameter $\alpha$ measures the sensibility of $P$ to current situations, and $Q_{t}:=\frac{1}{N}\sum_{i=1}^{N}v_{t}^{i}$ is the average trading rate at time $t$. The sticky price model well captures the persistent effects of the past price conditions on the current prices, reflecting realistic frictions in electricity markets.
@@
respectively. The parameter $\zeta$ corresponds to the preferred final storage, and $\kappa$ is the preferred instantaneous storage. We make the assumption that $\eta>0,\ \gamma\geq 0,\ c>0,\ \kappa\in\mathbb{R},\ \text{and}\ \zeta\in\mathbb{R}$ are constants. Compared with the conventional LQG framework, this cost function includes an additional $vP$ term. The introduction of this term weakens the coercivity condition, and increases the difficulty of estimating $\epsilon_{N}$ and deriving the optimal strategies.
@@
A set of admissive strategies $\hat{\boldsymbol{v}}=\{\hat{v}^{1},\cdots,\hat{v}^{N}\}\in\Pi_{i=1}^{N}\mathcal{A}^{i}$ is an $\epsilon$-Nash equilibrium if
$J_{i}(\hat{v}^{i},\hat{\boldsymbol{v}}^{-i})-\epsilon_{N}\leq\inf_{v^{i}\in\mathcal{A}_{i}}J(v^{i},\hat{\boldsymbol{v}}^{-i})\leq J_{i}(\hat{v}^{i},\hat{\boldsymbol{v}}^{-i}),\quad\forall i=1,\ldots,N,$ (2.4)
@@
Under the second case (namely, the cooperate game case), instead of controlling the electricity independently, the agents cooperate with each other, and aim to achieve the best collective outcome. In this article, the collective outcome represents the average value of $J_{i}(v^{i},\boldsymbol{v}^{-i})$. It reads
@@
$|J_{soc}(\check{v}^{1},\check{v}^{2},\cdots,\check{v}^{N})-\inf_{v^{i}\in\mathcal{A}_{i},1\leq i\leq N}J_{soc}(v^{1},v^{2},\cdots,v^{N})|\leq\epsilon_{N},$ (2.5)
where $\epsilon_{N}$ goes to 0 as $N\to\infty$.
Here, the restriction for $\epsilon_{N}$ reflects that the expected social cost under the proposed strategies converges to that of the social optimal strategies as $N$ goes to $\infty$.
@@
3 Non-cooperate Game
In this section, we aim to find an $\epsilon$-Nash equilibrium for this system. We will use a mean-field approximation approach, in which we construct $C[0,T]$ functions to approximate certain processes. Our approach can be outlined in the following steps.
@@ 3 Non-cooperate Game
    The first step in constructing a mean-field approximation is to approximate $Q$ with a continuous function $\bar{Q}=(\bar{Q}(t))_{t\in[0,T]}$. It follows from (2.2) that the approximate price function $(\bar{P}(t))_{t\in[0,T]}$ satisfies
@@
    For each agent $i$, solve the auxiliary control problem (3.2) to find the optimal control $\hat{v}^{i}$. We calculate the average of the corresponding optimal controls and remove the noise term to obtain a function which measures the expectation of the average of the optimal controls. By requiring that this function equal $\bar{Q}$, we determine the exact form of $\bar{Q}$.
@@
The rationale for proposing the first two steps is primarily based on the following reasons. Firstly, when the population enlarges to $\infty$, the representative agent has no influence on $Q$, and therefore has no influence on $P$, as but one agent amid a continuum. From its perspective, the process $P$ should be evolved in its cost as a deterministic function. Secondly, the optimization problem for (2.3) becomes much easier to solve once we treat $P$ as a fixed continuous function, instead of a stochastic process. For each $\bar{Q}$, a unique process $\bar{P}$ is determined by (3.1). Each agent determines its best response to minimize (3.2). By calculating the expectation of the average of the optimal controls, we derive a function determined by $\bar{Q}$. If the expectation of this function is equal to $\bar{Q}$, a closed-loop is formed. Within this loop, the strategy adopted by each agent yields outcomes that closely approximate those of the actual optimal strategy as the number of agents increases.
For the given function $\bar{Q}\in C[0,T]$, we derive the approximate price process $\bar{P}$ by (3.1). We propose the value function of the auxiliary control problem (3.2),
$K_{i}(t,x)=\inf_{v\in\mathcal{B}_{i}}\mathbb{E}\left[\int_{t}^{T}(L(X_{s},v_{s})+\bar{P}(s)v_{s})ds+\Psi(X_{T})\right].$
where $t$ represents the starting time, $X$ satisfies
$dX_{s}=v_{s}ds+\sigma_{i}dW_{s},\quad\forall t\leq s\leq T;\quad X_{t}=x,$ (3.3)
and the admissible set for $v$ is denoted by
[…]
By dynamic programming principle, the HJB equation satisfies
$\frac{\partial{K}_{i}(t,x)}{\partial t}+\inf_{v^{i}\in\mathcal{B}_{i}}\left\{v_{t}^{i}\frac{\partial{K}_{i}(t,x)}{\partial x}+\frac{c}{2}({v_{t}^{i}})^{2}+v_{t}^{i}\bar{P}(t)\right\}+\frac{\sigma_{i}^{2}}{2}\frac{\partial^{2}{K}_{i}(t,x)}{\partial x^{2}}+\frac{\eta}{2}(x-\kappa)^{2}=0,$ (3.4)
along with the terminal condition $K_{i}(T,x)=\frac{\gamma}{2}(x-\zeta)^{2}$. We make the ansatz ${K}_{i}(t,x)=a_{i}(t)x^{2}+B_{i}(t)x+F_{i}(t)$ yields. Furthermore, the optimal feedback control achieving the minimum is given by
$\hat{v}_{t}^{i}=-\frac{\bar{P}(t)+2a_{i}(t)x+B_{i}(t)}{c},\quad\forall t\in[0,T].$ (3.5)
By substituting (3.5) into (3.4), we obtain that
$\qquad\qquad\qquad a^{\prime}_{i}(t)+\frac{\eta}{2}-\frac{(2a_{i}(t))^{2}}{2c}=0,$ $a_{i}(T)=\frac{\gamma}{2},$ (3.6)
$\qquad\qquad\qquad B^{\prime}_{i}(t)-\eta\kappa-\frac{2a_{i}(t)(\bar{P}(t)+B_{i}(t))}{c}=0,$ $B_{i}(T)=-\gamma\zeta,$ (3.7)
$\qquad\qquad\qquad F^{\prime}_{i}(t)-\frac{(\bar{P}(t)+B_{i}(t))^{2}}{2c}+\frac{\eta\kappa^{2}}{2}+\sigma_{i}^{2}a(t)=0,$ $F_{i}(T)=\frac{\gamma\zeta^{2}}{2}.$ (3.8)
We will subsequently verify that (3.6)–(3.8) indeed admit a solution. To this end, we introduce section 3.
@@
The solution to (3.6), (3.7) and (3.8) exists and is unique.
Remark 2.
Throughout this paper, unless stated otherwise, we seek classical solutions to the equations under consideration. Since the solution to (3.6) and (3.7) is uniform over $i$, so in the following, we abbreviate $a_{i}(t),B_{i}(t)$ as $a(t),B(t)$, respectively.
The proofs of the Propositions in this paper are collected to Appendix B.
Next, we present the verification theorem for the best-response control that minimizes (3.2).
@@
$\hat{v}_{t}^{i}=-\frac{\bar{P}(t)+2a(t){X}_{t}^{i}+B(t)}{c},\quad\forall t\in[0,T],$
where $a(t),B(t)$ are given by (3.6) and (3.7). Then, the SDE
@@
admits a unique strong solution, denoted by $\hat{X}_{t}^{i}$, and $\hat{v}^{i}$ is an optimal Markovian control for minimizing (3.2).